OpenStudy (anonymous):

The sum of two consecutive odd integers is 56. What is the smaller number?

OpenStudy (anonymous):

@ikatouni @undeadknight26

OpenStudy (cwrw238):

n + n + 2 = 56

OpenStudy (anonymous):

do u want the choices??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 56?? @ikatouni

OpenStudy (cwrw238):

no solve the equation for n what do u get?

OpenStudy (anonymous):

ok!! thanks!!

OpenStudy (anonymous):

The sum of the two is n + (n+2) = 56. Now go solve it

OpenStudy (cwrw238):

2n + 2 = 56 can you find what n is?

OpenStudy (anonymous):

@faylynn find the ,issing numbed

undeadknight26 (undeadknight26):

hmm may i offer a easier solution? take 56/2 = 28 28 + 28 = 56 -1 +2 27 + 29 = 56....does this help you any?

OpenStudy (anonymous):

@faylynn if you add 2 to the smaller odd number you end with the larger odd number y = x + 2 sum: `````` x + y = 56 substitute with y = x + 2 into x + y = 56 x + y = 56 x + (x +2) = 56 2x + 2 = 56 2x = 54 x = 27 y = x + 2 y = 27 + 2 y = 29 smaller number = 27

OpenStudy (anonymous):

ok!! hang on!!

OpenStudy (anonymous):

@faylynn ur answer is 27

undeadknight26 (undeadknight26):

You guys got it! Now let me go to taco bell before breakfast is off ;-;

OpenStudy (anonymous):

thanks!! both of u!! i cant give u both a medal!! but can fan!! lolz

OpenStudy (anonymous):

thank u all!! all of u r lifesavers!!